How many grams of HCl is needed to neutralize 2.0 L of a 0.1M of CaCO3?
Answer #1

Balanced chemical equation is:

CaCO3 + 2 HCl ---> CaCl2 + CO2 + H2O

lets calculate the mol of CaCO3

volume , V = 2 L

use:

number of mol,

n = Molarity * Volume

= 0.1*2

= 0.2 mol

According to balanced equation

mol of HCl reacted = (2/1)* moles of CaCO3

= (2/1)*0.2

= 0.4 mol

This is number of moles of HCl

Molar mass of HCl,

MM = 1*MM(H) + 1*MM(Cl)

= 1*1.008 + 1*35.45

= 36.458 g/mol

use:

mass of HCl,

m = number of mol * molar mass

= 0.4 mol * 36.46 g/mol

= 14.58 g

Answer: 14.6 g